# BetterV: Controlled Verilog Generation with Discriminative Guidance

### v1

@@ Abstract
Due to the growing complexity of modern Integrated Circuits (ICs), there is a need for automated circuit design methods. Recent years have seen rising research in hardware design language generation to facilitate the design process. In this work, we propose a Verilog generation framework, BetterV, which fine-tunes the large language models (LLMs) on processed domain-specific datasets and incorporates generative discriminators for guidance on particular design demands. The Verilog modules are collected, filtered and processed from internet to form a clean and abundant dataset. Instruct-tuning methods are specially designed to fine-tuned the LLMs to understand the knowledge about Verilog. Furthermore, data are augmented to enrich the training set and also used to train a generative discriminator on particular downstream task, which leads a guidance for the LLMs to optimize the Verilog implementation. BetterV has the ability to generate syntactically and functionally correct Verilog, which can outperform GPT-4 on the VerilogEval-machine benchmark. With the help of task-specific generative discriminator, BetterV can achieve remarkable improvement on various electronic design automation (EDA) downstream tasks, including the netlist node reduction for synthesis and verification runtime reduction with Boolean Satisfiability (SAT) solving.
@@ 1 Introduction
Large language models (LLMs) are leading the world because of their strong capability on generating and understanding natural language at a massive scale, which makes their potential applications and benefits for various domains and tasks. On the field of coding, LLMs can act as a seasoned assistant for developers to give advise on programming, find and fix bugs on the code and even generate the whole code with descriptions (Chen et al., 2021; Nijkamp et al., 2022, 2023; Roziere et al., 2023).
@@
Electronic design automation (EDA) is a set of softwares and services for designing integrated circuits (ICs), which work together in design flow to conceive and analyze the circuit designs. The slowing down of the Moore’s law puts an increasing pressure on EDA, and there is an emergent need to further improve and automate the design flow. Therefore, it is expected to incorporate LLMs into EDA flow. LLMs on EDA have already achieved remarkable success on various tasks (Liu et al., 2023a; He et al., 2023; R et al., 2023).
@@
Hardware design languages (HDLs), such as Verilog and VHDL, describe the hardware design at the very beginning of the design flow, which play an important role in the EDA flow and have strong influence on the following stages. However, writing the HDL is time-consuming and bug-prone, which makes it more expensive for today’s complex ICs. Therefore, it is promising to utilize the LLMs to automatically generate the desired HDL. Recently, there have been several works focus on Verilog generation (Thakur et al., 2023; Dehaerne et al., 2023). However, these works only focus on refining LLMs with selected datasets, neglecting the direct integration of syntactic or functional correctness in the model’s development. Moreover, they insufficiently address pivotal downstream tasks in the EDA process, which should be crucial evaluations for the generated Verilog.
@@
Despite that we expect that LLMs can play an important role in Verilog generation, there still are some challenges that we need to figure out. Firstly, the complex and strict requirements of hardware designs restrain the LLMs to learn and understand the knowledge related to the Verilog. Secondly, there are limited Verilog resources available in the world, which always leads to the problems of overfitting and data bias for LLMs fine-tuning. Moreover, taking the complicated and various downstream tasks into consideration further makes it a challenge, since it is not practical to fine-tune the LLMs for each downstream task.
@@
The controllable text generation contains techniques to train extra discriminators to guide the LLMs on desired directions (Zhang et al., 2023; Holtzman et al., 2018; Scialom et al., 2020). However, these works only pay attention to control the natural language text representation in their development. In this work, we try tapping into the potential to use this technique on optimization tasks, which is more difficult and different from simply treating the text representation.
In this work, we propose a framework, BetterV, for Verilog generation, by instruct-tuning the LLMs on our processed datasets and incorporating generative discriminators to optimize the Verilog implementation on various downstream tasks. It successfully teaches the LLMs to understand domain-specific knowledge and can adapt to any Verilog related downstream tasks.
@@
  • Utilizing a fine-tuned 7B-parameter LLM, without the application of prompt-engineering strategies, BetterV demonstrates the capacity to generate syntactically and functionally correct Verilog, which surpasses GPT-4 when evaluated on the VerilogEval-machine benchmark.
  • BetterV offers a versatile solution for data augmentation, tailored to meet diverse specifications in Verilog implementation. This approach addresses the challenge posed by the scarcity of Verilog resources effectively.
@@ 2.1 LLMs for Verilog Generation
Large language models (LLMs) have shown remarkable performance on code generation with either training a model from the beginning or fine-tuning a pre-trained model (Chen et al., 2021; Nijkamp et al., 2022, 2023; Roziere et al., 2023). The success of LLMs on code generation arouses the interest of study of LLMs on hardware design. As a widely recognized hardware description language (HDL), the generation of Verilog using LLMs has garnered significant attention and undergone extensive exploration. Some studies pay attention to construct customized datasets for the fine-tuning of LLMs, such as (Thakur et al., 2023) and (Dehaerne et al., 2023), who collect Verilog from the internet and process the data prior to training. VerilogEval (Liu et al., 2023b) and RTLCoder (Liu et al., 2023c) further study the importance of problem descriptions and then generate various problem-answer pairs as dataset. RTLCoder also considers the quality feedback on different data by ranking them with scores during the training. Researchers also try utilizing the prompt-engineering techniques to enhance the generation ability, such as the self-planning used in RTLLM (Lu et al., 2023) and the Retrieval-Augmented generation (RAG) in RTLFixer (Tsai et al., 2023). VerilogEval (Liu et al., 2023b) and RTLLM (Lu et al., 2023) build benchmarks to evaluate the generated Verilog on their functional or syntactic correctness.
@@ 2.2 Discriminator-guided Controllable Generation
Controlling LLMs has been widely explored during recent years (Zhang et al., 2023). Class-conditional language models (CCLMs), such as CTRL (Keskar et al., 2019), aim at controlling the generation by appending a control code in the beginning of training sequences. Discriminator-guided controllable generation is an important technique used for controllable generation, which combines a discriminator with the generative LLMs. The constraints are modeled by calculating the conditional probability on each class and related to the next-token probabilities by Bayes rule. (Holtzman et al., 2018) and (Scialom et al., 2020) predict the labels by feeding each candidate next token into a discriminator and hence guiding the generation to desired directions. PPLM (Dathathri et al., 2020) further employs a forward and backward process with the gradient from a discriminator to update the latent states of LLMs to guide the target generation. In order to reduce the cost of using discriminator for each possible next tokens, GeDi (Krause et al., 2020) contrastively trains the CCLMs as generative discriminators to guide the generation during the decoding. DEXPERTS (Liu et al., 2021) attempts to further improve the performance by incorporating expert and anti-expert during the decoding.
Figure 1: The overview of BetterV.
@@ 3.2 Instruct-Tuning Data-Processing
Inspired by (Dehaerne et al., 2023), we collect open-source repositories in GitHub that containing Verilog or SystemVerilog. At the same time, the repository licenses are also checked to permit modification and distribution. We also filter out the auto-generated files that are mostly repetitive and the non-permissive files that contradict the licenses. As in (Dehaerne et al., 2023), the files with too much or less lines are filtered. These files are then processed by extracting the Verilog modules and functions with regular expressions. We further analyze the extracted contents by measuring the token number after encoding them with the tokenizer and then remove the contents exceeding a pre-defined maximum token number. The reason is because the contents that exceed the maximum token number will be truncated by the tokenizer during training and hence it cannot provide enough meaningful information with incomplete contents.
In order to further support our instruct-tuning tasks, we implement two kinds of processing on the collected contents. The first one is that we use a V2C tool to translate Verilog into C (Mukherjee et al., 2016), then the hardware design information represented in Verilog is also transferred into the C program. The dataset is then appended by the translated C programs, which are able to be translated from the Verilog and also don’t exceed the maximum token number. Secondly, we split the Verilog modules and functions into their definition (with the module header, input and output definition) and body.
@@ 3.3 Domain-specific Instruct-Tuning
To solve this issue, we propose domain-specific instruct-tuning, where examples are shown in Figure 2. We make use of the abundant knowledge already learnt by LMMs on general code programs. Then we construct an instruction that asks the LLMs to translate the Verilog into C and also translate from C to Verilog, by which LLMs can learn to understand Verilog from the implementation of C. With the help of V2C tool, the dataset is easy to be obtained as described in last section. Besides this, in order to improve the performance of LLMs to follow the Verilog generation instruction, the instruction on Verilog autocompletion is constructed, where the module definition is in the instruction and the completed module is the answer. It should be noted that we don’t add extra natural language description in instruct tuning since the implementation of C program already acts as functional description to teach the LLMs.
To be noticed, the LLMs follow the an auto-regressive manner, where a LLM with parameter $\theta$ predicts the probability of a sequence $x=\{x_{1},...,x_{n}\}$ with both instruction and answer and factorizes it using the chain rule of probability as follows (Keskar et al., 2019):
@@ 3.5 Generative Discriminator
For generative discriminator, we have binary cases to represent opposite attributes, i.e. a control code $c$ and an anti-control code $\bar{c}$ (Krause et al., 2020), and the inputs are respectively conditioned by them as $p_{\theta}(x|c)$ and $p_{\theta}(x|\bar{c})$ to guide the LLMs on $p_{LLM}(x)$. In our domain-specific scenarios, the preceding codes indicate which kind of Verilog attribute we desired or undesired. As described in Section 3.4, we expect the generative discriminator can generate correct and better Verilog implementation, and their labels are corresponding to the conditioned attributes.
@@
With this generative discriminator training pipeline and the equipped downstream task-specific augmented data, we can model any optimization scenario related to Verilog implementation and train a distinct discriminator to guide towards our desired directions. Moreover, with the updated Verilog as new reference, we can iteratively update the discriminator to continually optimize the Verilog.
@@ 4 Experiments
In this section, we first introduce the experimental setting as well as the evaluation metrics in Section 4.1. Then in Section 4.2 we present our results on functional correctness and compared with other methods, followed by further results with different downstream tasks incorporated with the discriminator (including Syntactic Correctness in Section 4.3, Synthesis Nodes Reduction in Section 4.4.1 and Verification Runtime Reduction in Section 4.4.2).
@@ 4.1 Experimental Setting
We choose the CodeLlama-7B-Instruct (Roziere et al., 2023) model as our pre-trained generative LLM, and the TinyLlama (Zhang et al., 2024) model as our pre-trained generative discriminator. BetterV is trained with the help of DeepSpeed ZeRO (Rajbhandari et al., 2020) and LoRA (Hu et al., 2021) on the DeepSpeed-Chat (Yao et al., 2023). The experiments are conducted on a machine with two NVIDIA Tesla V100S PCIe 32 GB graphics cards with CUDA driver 11.4.
We employ the VerilogEval (Liu et al., 2023b), which comprises various problems with either machine-generated or human-crafted, as our evaluation benchmark. Following VerilogEval, we measure the Verilog functional correctness with simulation through the pass@k metric with unbiased estimator:
@@
Due to the Plug-and-Play nature of the discriminator, BetterV can work on various downstream tasks. For other downstream tasks besides the functional correctness, we can still utilize the pass@k metric and the problems in VerilogEval but the correctness is measured by different tools in $\mathsf{Yosys}$ (Wolf, ) according to different downstream task. In our experiments, we sample n = 20 code completions per problem for each downstream task and measuring pass@k with k = 1,5,10.
For all the models, we employ the Adam optimizer (Kingma & Ba, 2014) with $\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.95 and the cosine learning rate decay (Loshchilov & Hutter, 2016) to schedule our learning rate. For the generative 7B LLMs fine-tuning process, we train it for 4 epochs using an initial learning rate of 9.65e-6 with a batch size of 4. For the generative discriminator, we train it for 3 epochs using an initial learning rate of 9.65e-6 with a batch size of 8. The LoRA dimension for both LLMs and discriminator is set as 128. For different downstream tasks they have different sensitivity for the incorporation of generative discriminators, hence the setting of $\lambda$, $w$, $\rho$ and $\tau$ are different and task-sensitive. We initialize the value of learnable scale $\alpha$ as $1$ and bias $b$ as 00.
4.2 Functional Correctness
For functional correctness, the target is to correctly complete the Verilog module given the problem description and module definition. The performance of BetterV is compared with GPT-3.5, GPT-4, CodeLlama-7B-Instruct (Roziere et al., 2023), VerilogEval (Liu et al., 2023b), ChipNeMo (Liu et al., 2023a) and RTLCoder (Liu et al., 2023c). As shown in Table 1, the results demonstrate that our BetterV have achieved the state-of-the-art performance on VerilogEval-machine, where the pass@1, pass@5 and pass@10 outperform GPT-4 by $4.2$, $4.8$ and $5.6$, respectively. BetterV can also perform well on VerilogEval-human and outperform all the baselines except the GPT-4. The reason that BetterV doesn’t outperform GPT-4 on VerilogEval-human may because that lots of the human-crafted problem descriptions in VerilogEval-human are made by directly converting non-text descriptions into text-only structure, which are always brand-new knowledge for both pre-trained model and our fine-tuned model. It should be noted that the pre-trained model CodeLlama-7B-Instruct has the lowest performance, but after instruct-tuning and the guidance of discriminator a huge performance are improved.
@@ 4.2 Functional Correctness
ChipNeMo 43.4 - - 22.4 - -
RTLCoder 62.5 - - 36.7 - -
BetterV 64.2 75.4 79.1 40.9 50.0 53.3
We also show that the generative discriminator has the ability to further enhance the performance after the guidance. In the task of functional correctness, the labels of discriminator are made by checking whether the generated modules are functional equivalence with the reference module. The equivalence checking is done by the $\mathsf{eqy}$ tool in $\mathsf{Yosys}$. “BetterV-base” means the base model of BetterV that fine-tuend after instruct-tuning without discriminator. With the problems in VerilogEval-human, the results in Table 2 illustrate that the guidance from the discriminator can not only enhance the capability on our fine-tuned LLMs, i.e. BetterV-base, but also on the original pre-trained LLMs, i.e. CodeLlama. This observation indicates that our trained discriminator can be employed to any LLM, only if they have the same vocabulary size.
Table 2: Improvements of functional correctness with discriminator.
Model pass@1 pass@5 pass@10
CodeLlama 18.2 22.7 24.3
CodeLlama + Dis 20.3 24.1 24.7
BetterV-base 40.0 49.5 53.0
BetterV 40.9 50.0 53.3
4.3 Syntactic Correctness
Table 3: Improvements of syntactic correctness with discriminator.
Model pass@1 pass@5 pass@10
CodeLlama 41.5 50.8 53.8
CodeLlama + Dis 49.1 58.4 61.1
BetterV-base 82.6 97.6 99.2
BetterV 87.1 98.2 99.3

### v2

@@ Abstract
Due to the growing complexity of modern Integrated Circuits (ICs), there is a need for automated circuit design methods. Recent years have seen rising research in hardware design language generation to facilitate the design process. In this work, we propose a Verilog generation framework, BetterV, which fine-tunes the large language models (LLMs) on processed domain-specific datasets and incorporates generative discriminators for guidance on particular design demands. The Verilog modules are collected, filtered and processed from internet to form a clean and abundant dataset. Instruct-tuning methods are specially designed to fine-tune the LLMs to understand the knowledge about Verilog. Furthermore, data are augmented to enrich the training set and also used to train a generative discriminator on particular downstream task, which leads a guidance for the LLMs to optimize the Verilog implementation. BetterV has the ability to generate syntactically and functionally correct Verilog, which can outperform GPT-4 on the VerilogEval benchmark. With the help of task-specific generative discriminator, BetterV can achieve remarkable improvement on various electronic design automation (EDA) downstream tasks, including the netlist node reduction for synthesis and verification runtime reduction with Boolean Satisfiability (SAT) solving.
@@ 1 Introduction
Large language models (LLMs) are leading the world because of their strong capability on generating and understanding natural language at a massive scale, which makes their potential applications and benefits for various domains and tasks. On the field of coding, LLMs can act as a seasoned assistant for developers to give advise on programming, find and fix bugs on the code and even generate the whole code with descriptions [2, 20].
@@
Electronic design automation (EDA) is a set of softwares and services for designing integrated circuits (ICs), which work together in design flow to conceive and analyze the circuit designs. The slowing down of the Moore’s law puts an increasing pressure on EDA, and there is an emergent need to further improve and automate the design flow. Therefore, it is expected to incorporate LLMs into EDA flow. LLMs on EDA have already achieved remarkable success on various tasks [13, 6].
@@
Hardware design languages (HDLs), such as Verilog and VHDL, describe the hardware design at the very beginning of the design flow, which play an important role in the EDA flow and have strong influence on the following stages. However, writing the HDL is time-consuming and bug-prone, which makes it more expensive for today’s complex ICs. Therefore, it is promising to utilize the LLMs to automatically generate the desired HDL. Recently, there have been several works focus on Verilog generation [24, 4]. However, these works only focus on refining LLMs with selected datasets, neglecting the direct integration of syntactic or functional correctness in the model’s development. Moreover, they insufficiently address pivotal downstream tasks in the EDA process, which should be crucial evaluations for the generated Verilog.
@@
Despite that we expect that LLMs can play an important role in Verilog generation, there still are some challenges that we need to figure out. Firstly, the complex and strict requirements of hardware designs restrain the LLMs to learn and understand the knowledge related to the Verilog. Secondly, there are limited Verilog resources available in the world, which always leads to the problems of overfitting and data bias for LLMs fine-tuning. Moreover, taking the complicated and various downstream tasks into consideration further makes it a challenge, since in the industrial production the optimize of downstream targets from Verilog design to physical implementation is various and can’t be accomplished in an action, which always need many loops in the industrial production. Therefore, it’s difficult for LLMs to look ahead and it’s also not practical to fine-tune the LLMs for each downstream task.
@@
The controllable text generation contains techniques to train extra discriminators to guide the LLMs on desired directions [23]. However, these works only pay attention to control the natural language text representation in their development. In this work, we try tapping into the potential to use this technique on optimization tasks, which is more difficult and different from simply treating the text representation.
In this work, we propose a framework, BetterV, for Verilog generation, by instruct-tuning the LLMs on our processed datasets and incorporating generative discriminators to optimize the Verilog implementation on various downstream tasks. We utilize the alignment between Verilog and the C program to help the LLMs understand the Verilog effectively. At the same time, data augmentation is proposed to solve the Verilog data scarce issue. What’s more, we consider that there are many implementations of Verilog that always have different performance on PPA (Power, Performance, Area) or verification runtime. Therefore, the problems in downstream tasks are transformed to how to optimize the Verilog implementation. The generative discriminator is then employed to guide the LLMs to generate or modify Verilog implementation directly from natural language and be as friendly to do downstream tasks as possible at the same time, thereby effectively reducing the number of iterations on industrial production. With the above techniques, BetterV successfully teaches the LLMs to understand domain-specific knowledge and can adapt to any Verilog related downstream tasks.
@@
  • Utilizing fine-tuned 6.7B/7B-parameter LLMs, without the application of prompt-engineering strategies, BetterV demonstrates the capacity to generate syntactically and functionally correct Verilog, which surpasses GPT-4 when evaluated on the VerilogEval benchmark.
  • BetterV offers a versatile solution for data augmentation, tailored to meet diverse specifications in Verilog implementation. This approach addresses the issue posed by the scarcity of Verilog resources effectively.
@@ 2.1 LLMs for Verilog Generation
Large language models (LLMs) have shown remarkable performance on code generation with either training a model from the beginning or fine-tuning a pre-trained model [20, 22]. The success of LLMs on code generation arouses the interest of study of LLMs on hardware design. As a widely recognized hardware description language (HDL), the generation of Verilog using LLMs has garnered significant attention and undergone extensive exploration. Some studies pay attention to construct customized datasets for the fine-tuning of LLMs, such as [24] and [4], who collect Verilog from the internet and process the data prior to training. VerilogEval [14] and RTLCoder [15] further study the importance of problem descriptions and then generate various problem-answer pairs as dataset. RTLCoder also considers the quality feedback on different data by ranking them with scores during the training. Researchers also try utilizing the prompt-engineering techniques to enhance the generation ability, such as the self-planning used in RTLLM [17]. VerilogEval [14] and RTLLM [17] build benchmarks to evaluate the generated Verilog on their functional or syntactic correctness.
@@ 2.2 Discriminator-guided Controllable Generation
Controlling LLMs has been widely explored during recent years. Class-conditional language models (CCLMs), such as CTRL [9], aim at controlling the generation by appending a control code in the beginning of training sequences. Discriminator-guided controllable generation is an important technique used for controllable generation, which combines a discriminator with the generative LLMs. The constraints are modeled by calculating the conditional probability on each class and related to the next-token probabilities by Bayes rule. [7] and [23] predict the labels by feeding each candidate next token into a discriminator and hence guiding the generation to desired directions. PPLM [3] further employs a forward and backward process with the gradient from a discriminator to update the latent states of LLMs to guide the target generation. In order to reduce the cost of using discriminator for each possible next tokens, GeDi [11] contrastively trains the CCLMs as generative discriminators to guide the generation during the decoding. DEXPERTS [12] attempts to further improve the performance by incorporating expert and anti-expert during the decoding.
@@ 3.2 Instruct-Tuning Data-Processing
Inspired by [4], we collect open-source repositories in GitHub that contain Verilog or SystemVerilog. At the same time, the repository licenses are also checked to permit modification and distribution. We also filter out the auto-generated files that are mostly repetitive and the non-permissive files that contradict the licenses. As in [4], the files with too much or less lines are filtered. These files are then processed by extracting the Verilog modules and functions with regular expressions. We further analyze the extracted contents by measuring the token number after encoding them with the tokenizer and then remove the contents exceeding a pre-defined maximum token number. The reason is because the contents that exceed the maximum token number will be truncated by the tokenizer during training and hence it cannot provide enough meaningful information with incomplete contents. Finally, we filtered out the Verilog modules that have syntactic error.
In order to further support our instruct-tuning tasks, we implement two kinds of processing on the collected contents. The first one is that we use a V2C tool to translate Verilog into C [19], then the hardware design information represented in Verilog is also transferred into the C program. The dataset is then appended by the translated C programs, which are able to be translated from the Verilog and also don’t exceed the maximum token number. Secondly, we split the Verilog modules and functions into their definition (with the module header, input and output definition) and body.
@@ 3.3 Domain-specific Instruct-Tuning
To solve this issue, we propose domain-specific instruct-tuning, where examples are shown in Figure 2. Firstly, we introduce a novel alignment method that maps Verilog code to C, facilitating LLM’s understanding of Verilog. We make use of the abundant knowledge already learnt by LLMs on general code program, i.e. the C program. Then we construct an instruction that asks the LLMs to translate the Verilog into C and also translate from C to Verilog, by which LLMs can learn to understand Verilog from the implementation of C. With the help of V2C tool, the dataset is easy to be obtained as described in last section. Besides this, in order to improve the performance of LLMs to follow the Verilog generation instruction, the instruction on Verilog autocompletion is constructed, where the module definition is in the instruction and the completed module is the answer. It should be noted that we don’t add extra natural language description in instruct tuning since the implementation of C program already acts as functional description to teach the LLMs.
To be noticed, the LLMs follow the an auto-regressive manner, where a LLM with parameter $\theta$ predicts the probability of a sequence $x=\{x_{1},...,x_{n}\}$ with both instruction and answer and factorizes it using the chain rule of probability as follows [9]:
@@ 3.5 Generative Discriminator
For generative discriminator, we have binary cases to represent opposite attributes, i.e. a control code $c$ and an anti-control code $\bar{c}$ [11], and the inputs are respectively conditioned by them as $p_{\theta}(x|c)$ and $p_{\theta}(x|\bar{c})$ to guide the LLMs on $p_{LLM}(x)$. In our domain-specific scenarios, the preceding codes indicate which kind of Verilog attribute we desired or undesired. As described in Section 3.4, we expect the generative discriminator can generate correct and better Verilog implementation, and their labels are corresponding to the conditioned attributes.
@@
With this generative discriminator training pipeline and the equipped downstream task-specific augmented data, we can model any optimization scenario related to Verilog implementation and train a distinct discriminator to guide towards our desired directions. At the same time, the loops in the industrial production and the manual cost can be reduced effectively by jointly considering the corresponding downstream constraints at the phase of Verilog generation.
@@ 4 Experiments
In this section, we first introduce the experimental setting as well as the evaluation metrics in Section 4.1. Then in Section 4.2 we present our results on functional correctness and compared with other methods, followed by further results with different downstream tasks incorporated with the discriminator (including Synthesis Nodes Reduction in Section 4.3.1 Verification Runtime Reduction in Section 4.3.2) and ablation study in Section 4.4.
@@ 4.1 Experimental Setting
We fine-tune the CodeLlama-7B-Instruct [22] and DeepSeek-Coder-6.7b-Instruct [5] as generative LLMs, and fine-tune TinyLlama [27] and DeepSeek-Coder-1.3b-Instruct [5] as generative discriminator, respectively. For CodeQwen1.5-7B-Chat [1], we can only fine-tune it since there is no suitable smaller size version of it. BetterV is trained with the help of DeepSpeed ZeRO [21] and LoRA [8] on the DeepSpeed-Chat [26]. The experiments are conducted on a machine with two NVIDIA Tesla V100S PCIe 32 GB graphics cards with CUDA driver 11.4.
We employ the VerilogEval [14], which comprises various problems with either machine-generated or human-crafted, as our evaluation benchmark. Following VerilogEval, we measure the Verilog functional correctness with simulation through the pass@k metric with unbiased estimator:
@@
Due to the Plug-and-Play nature of the discriminator, BetterV can work on various downstream tasks. For other downstream tasks besides the functional correctness, we can still utilize the pass@k metric and the problems in VerilogEval but the correctness is measured by different tools in $\mathsf{Yosys}$ [25] according to different downstream task. In our experiments, we sample n = 20 code completions per problem for each downstream task and measuring pass@k with k = 1,5,10.
For all the models, we employ the Adam optimizer [10] with $\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.95 and the cosine learning rate decay [16] to schedule our learning rate. For the generative LLMs fine-tuning process, we train it for 4 epochs using an initial learning rate of 9.65e-6 with a batch size of 4. For the generative discriminator, we train it for 3 epochs using an initial learning rate of 9.65e-6 with a batch size of 8. The LoRA dimension for both LLMs and discriminator is set as 128. For different downstream tasks they have different sensitivity for the incorporation of generative discriminators, hence the setting of $\lambda$, $w$, $\rho$ and $\tau$ are different and task-sensitive. We initialize the value of learnable scale $\alpha$ as $1$ and bias $b$ as 00.
@@
RTLCoder-Mistral 62.5 72.2 76.6 36.7 45.5 49.2
RTLCoder-DeepSeek 61.2 76.5 81.8 41.6 50.1 53.4
BetterV-CodeLlama 64.2 75.4 79.1 40.9 50.0 53.3
BetterV-DeepSeek 67.8 79.1 84.0 45.9 53.3 57.6
BetterV-CodeQwen 68.1 79.4 84.5 46.1 53.7 58.2
4.2 Functional Correctness
For functional correctness, the target is to correctly complete the Verilog module given the problem description and module definition. The performance of BetterV is compared with GPT-3.5, GPT-4, CodeLlama-7B-Instruct [22], DeepSeek-Coder-6.7b-Instruct [5], CodeQwen1.5-7B-Chat [1], ChipNeMo [13], Thakur et al. [24], VerilogEval [14] and RTLCoder [15]. As shown in Table 1, the results demonstrate that our BetterV-CodeQwen have achieved the state-of-the-art performance on VerilogEval, where the pass@1 on VerilogEval-machine and VerilogEval-human outperforms GPT-4 by $8.1$ and $2.6$, respectively. It should be noted that the pre-trained model CodeLlama-7B-Instruct has the lowest performance, but after instruct-tuning and the guidance of discriminator a huge performance is improved.